# Black Hole from Entropy Maximization

### v4

@@ IV.2 Determination of the interior metric $g_{\mu\nu}^{*}$
and use two self-consistencies: thermodynamics and semi-classicality. Because we have considered quanta consistent with thermodynamics, the pressures must be positive Landau_SM . This and (30) require $C>0$ and $k-4\geq-2$, that is, $k\geq 2$. Also, we are focusing only on semi-classical configurations satisfying (17), which means that the curvatures $\mathcal{R}(r)$ must be at most of $\mathcal{O}(r^{0})$ for $r\gg l_{p}$. This restricts the highest term, the third one of (31), such that $2k-4\leq 0$, that is, $k\leq 2$. Therefore, we can only have $k=2$. From dimensional analysis, we can set $C=\frac{1}{2\sigma\eta}$ with a $\mathcal{O}(1)$ dimensionless constant $\eta>0$.
@@ IV.2.2 No singularity
$R_{RicciScalar}=-\frac{2}{L^{2}},~{}~{}R_{\mu\nu}R^{\mu\nu}=\frac{2}{L^{4}},~{}~{}R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}=\frac{4}{L^{4}},$ (34)
@@
We next examine the small center region $0\leq r\lesssim\sqrt{n}l_{p}$, which cannot be described by the metric (32). The energy inside is estimated through (24) as $\frac{a_{*}(\sqrt{n}l_{p})}{2G}\sim\sqrt{n}m_{p}$, which is much smaller than $\frac{a_{0}}{2G}$, one expected in a classical case. Therefore, the center region cannot have a singularity appearing in the classical cases, and rather it should correspond to a small excitation of quantum-gravitational degrees of freedom like string (which can be confirmed only by a future development). In this paper, we assume that the center part is almost flat Y1 . [^13] The metric (32) becomes flat around $r\approx\sqrt{2\sigma}$: $ds^{2}|_{r\approx\sqrt{2\sigma}}\approx-e^{\frac{1}{\eta}+A_{0}}dt^{2}+dr^{2}+r^{2}d\Omega^{2}$, which is a flat metric by redefining time coordinate $t$. Also, we can consider the time evolution from formation to evaporation and see that the center region is kept flat, except for the final stage of the evaporation KY4 .
Thus, the non-perturbative solution (32) with (33) has no singularity. We will see in Sec.IV.3.1 that a quantum repulsive force is generated inside and plays a key role in the resolution of the singularity.
@@ IV.3 Semi-classical gravity condensate
where $K$ is a $\mathcal{O}(1)$ number to be fixed. This is an exponentially large redshift KMY . To see this effect, setting $r=R-\Delta r$ ($\Delta r\ll R$), (37) becomes $-g_{tt}^{*}(r)\sim e^{-\frac{R\Delta r}{\sigma\eta}}$. This indicates that due to the strong redshift, from the outside, a deep region with $\Delta r\gg\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}}\right)$ is almost frozen in time. As a result, the information carried by the excited quanta is kept inside for an exponentially long time KY2 , and the gravity condensate looks almost black. (We will discuss its phenomenological effect in Sec.VII.)
IV.3.1 Formation process
We discuss how the gravity condensate is formed by a physical process. See Fig.4.
@@ IV.3.1 Formation process
Figure 4: Adiabatic formation of the semi-classical gravity condensate in a heat bath of Hawking temperature.
@@
Suppose that a collection of many spherical radiations at Hawking temperature comes together slowly due to self-gravity. This corresponds to the adiabatic formation process in a heat bath and can be done operationally by slowly changing the temperature of the heat bath to match the Hawking temperature of the energy at each stage KY1 .
@@
We can use (1) and make a self-consistent model that describes this process KMY ; KY1 ; KY4 ; KY5 ; Y1 . We first note that in this formation process, the metric changes in time, and that generically, particle creation occurs in a time-dependent spacetime BD ; Barcelo . Then, we can solve approximately quantum fields propagating in this metric to show that particles at the Hawking temperature are created in each part of the radiation. Considering the time evolution of both the collapsing matter and the backreaction from the particle creation by (1), [^15] It is often believed that the backreaction from evaporation is negligible when considering the formation process of a black hole with mass $\frac{a_{0}}{2G}$ since the time scale of the evaporation $\Delta t\sim\frac{a_{0}^{3}}{l_{p}^{2}}$ is much longer than the time scale of the collapse $\Delta\tau\sim a_{0}$. However, this naive idea is wrong because $t$ and $\tau$ are the time coordinate at infinity and that of the comoving observer along the collapsing matter, respectively, and it does not make sense to compare such two time scales. In fact, one can examine the time evolution of the evaporation and collapse in a common time coordinate, and check that the both effects compete near $r=a_{0}$ and the dense structure is formed KMY ; KY4 . the gravity condensate will be formed with the maximum entropy (25). This result is consistent with the second law of thermodynamics because reversible processes generally lead to most typical configurations in thermodynamics.
In the heat bath with the fixed Hawking temperature $T_{H}=\frac{\hbar}{4\pi a_{0}}$, the incoming radiation from the bath and the outgoing one from the condensate are balanced, and the system stays stationary KY2 . When taken out of the bath into a vacuum, the condensate evaporates in a time scale $\Delta t\sim\frac{a_{0}^{3}}{\sigma}$ KMY ; KY1 .
In this formation process, we can see three interesting points. First, we can find the origin of the large tangential pressure in (IV.3). Let us imagine that self-gravity brings together a large number of excited modes representing radiation from the bath, and the condensate is gradually formed. As the curvature increases, vacuum fluctuations of other modes with various angular momenta are induced causing the pressure as a non-perturbative effect KY4 (see Appendix B for a review). This can also be understood through the 4D Weyl anomaly KY3 . Because of the pressure, the excited quanta are not concentrated in the center but are distributed throughout the interior. As a result, the curvature remains finite as in (34), the center has only small energy, and no singularity appears.
@@
Second, the adiabatic formation is consistent with Bekenstein’s idea Bekenstein . At each stage where the mass is $\frac{a_{0}^{\prime}}{2G}$, a photon emitted from the bath of $T_{H}^{\prime}=\frac{\hbar}{4\pi a_{0}^{\prime}}$ comes to the gravity condensate of size $R^{\prime}=a_{0}^{\prime}+\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}^{\prime}}\right)$. Since the wavelength $\sim a_{0}^{\prime}$ and the size $R^{\prime}$ are almost the same, the probability for the photon to enter the dense region is roughly one-half, leading to 1 bit of information Bekenstein . On the other hand, the photon approaches to $r=R^{\prime}$  [^16] The model in Ref.KMY ; KY4 shows explicitly that, as a result of the self-consistent time-evolution considering the point of footnote 15, the photon will naturally approach $r=a^{\prime}_{0}+\frac{2\sigma}{a^{\prime}_{0}}$ with emitting particles at temperature $T_{H}^{\prime}=\frac{\hbar}{4\pi a_{0}^{\prime}}$., and the blueshifted temperature becomes $T_{loc}(r=R^{\prime})=\frac{T_{H}^{\prime}}{\sqrt{1-\frac{a_{0}^{\prime}}{R^{\prime}}}}\sim\frac{m_{p}}{\sqrt{n}}\sim\epsilon_{max}$. Therefore, this photon has $\epsilon_{max}$ and the 1 bit of information about its existence. Repeating this process, the uniform dense structure like Fig.2 appears, and the entropy follows the area law (25). In this sense, Bekenstein’s idea is realized, including the interior description at the level of (1).
Third, the semi-classical condition (18) is satisfied as a result of the time evolution by (1). As shown above, each photon approaches $r=R^{\prime}=a_{0}^{\prime}+\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}^{\prime}}\right)$, which makes the right hand side of (18) into $\sqrt{g_{rr}(R^{\prime})}(R^{\prime}-a_{0}^{\prime})\sim\sqrt{n}l_{p}$. On the other hand, the photon has the energy $\epsilon_{max}$, which means that the left hand side of (18) becomes $\frac{\hbar}{\epsilon_{max}}\sim\sqrt{n}l_{p}$. Then, the condition (18) is saturated at each stage of the formation. Note here that the condition (18) leads to the entropy bound (21), and that the formation process is reversible thermodynamically. Thus, these may imply that the condition (18), and hence the bound (21), is a consequence of the second law of thermodynamics at the level of the semi-classical Einstein equation (1).
@@ V.1.1 Local temperature
We first find the local temperature $T_{loc}^{*}(r)$. We note that the excited quanta composing the part around $r$ of the gravity condensate with the metric (32) are in radially accelerated motion against gravity to stay there as in Fig.2. The required acceleration is
@@
which is $\sim\epsilon_{max}$ consistent with (10). Thus, the excited quanta behave typically like a local thermal state in the radial direction at the temperature (39). This result is kinematical and can be applied to any type of local degrees of freedom in the metric (32), due to the universality of Unruh effect. On the other hand, we have seen in Sec.IV.3.1 that (39) is obtained dynamically through the particle creation during the formation process (see Y1 for an explicit proof). Therefore, the kinematical and dynamical results coincide, and (39) is robust.
@@
Therefore, we can understand (39) as follows. Each part of the interior remains, due to the large redshift, at the temperature $T_{loc}(r)\sim\epsilon_{max}$ at the time of formation (see Sec.IV.3.1), and only during $\Delta t<\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ can the gravity condensate exist consistent with Tolman’s law. In this sense, the entropy-maximizing configuration is not in global equilibrium but in radially local one  [^17] This non-global equilibrium makes the difference from the result of Ref.Oppenheim , and is consistent with the self-consistent state $|\psi\rangle_{*}$ (mentioned above (33)). See Sec.VII for discussions about a self-gravitating equilibrium state.. This is consistent with the result that Tolman’s law does not lead to maximum entropy (see Appendix D). Note that the time scale $\sim e^{a_{0}^{2}/l_{p}^{2}}$ is much longer than the evaporation time scale $\sim\frac{a_{0}^{3}}{l_{p}^{2}}$, and we can discuss the configuration physically.
@@ V.2 Surface
We here determine the relation of the size $R$ and the total energy $M_{0}=\frac{a_{0}}{2G}$ from thermodynamics Y1 . Imagine that the gravity condensate is in equilibrium with a heat bath of temperature $T_{0}$ during $\Delta t<\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ (see Right of Fig.4). From thermodynamic relation $T_{0}dS=dM_{0}$ and (42), the equilibrium temperature is determined as the Hawking temperature GH :
@@ VI.1 Derivation of the Bousso bound
We show that the Bousso bound holds in our class of configurations. We first check that the entropy evaluated on a $t$-constant spacelike hypersurface $\Sigma_{t}$ inside the surface at $r=R$, which we have considered so far, agrees with one evaluated on a spherical ingoing null hypersurface $\Sigma_{L}$ starting from the surface at $r=R$ and converging at $r=0$ (a light sheet) Bousso1 . We suppose here that all configurations satisfy (1) and are regular. Applying $\nabla_{\mu}\tilde{s}^{\mu}=0$ to a finite spacetime region $V$ enclosed with $\Sigma_{t}$ and $\Sigma_{L}$ ($\partial V=\Sigma_{L}-\Sigma_{t}$ as in Fig.5)

### v5

@@ IV.2 Determination of the interior metric $g_{\mu\nu}^{*}$
and use two self-consistencies: thermodynamics and semi-classicality. Because we have supposed quanta consistent with local thermodynamics, the pressures must be positive Landau_SM . This and (30) require $C>0$ and $k-4\geq-2$, that is, $k\geq 2$. Also, we are now considering the semi-classical condition (17), which means that the curvatures $\mathcal{R}(r)$ must be at most of $\mathcal{O}(r^{0})$ for $r\gg l_{p}$. This restricts the highest term, the third one of (31), such that $2k-4\leq 0$, that is, $k\leq 2$. Therefore, we can only have $k=2$. From dimensional analysis, we can set $C=\frac{1}{2\sigma\eta}$ with a $\mathcal{O}(1)$ dimensionless constant $\eta>0$.
@@ IV.2.2 No singularity
$R_{\rm RicciScalar}=-\frac{2}{L^{2}},~{}~{}R_{\mu\nu}R^{\mu\nu}=\frac{2}{L^{4}},~{}~{}R_{\mu\nu\alpha\beta}R^{\mu\nu\alpha\beta}=\frac{4}{L^{4}},$ (34)
@@
We next examine the small center region $0\leq r\lesssim\sqrt{n}l_{p}$, which cannot be described by the metric (32). The energy inside is estimated through (24) as $\frac{a_{*}(\sqrt{n}l_{p})}{2G}\sim\sqrt{n}m_{p}$, which is much smaller than $\frac{a_{0}}{2G}$, one expected in a classical case. Therefore, the center region cannot have a singularity appearing in the classical cases, and rather it should correspond to a small excitation of quantum-gravitational degrees of freedom like string (which can be confirmed only by a future development). In this paper, we assume that the center part is almost flat. [^13] The metric (32) becomes flat around $r\approx\sqrt{2\sigma}$: $ds^{2}|_{r\approx\sqrt{2\sigma}}\approx-e^{\frac{1}{\eta}+A_{0}}dt^{2}+dr^{2}+r^{2}d\Omega^{2}$, which is a flat metric by redefining time coordinate $t$. Also, we can consider the time evolution from formation to evaporation and see that the center region is kept flat, except for the final stage of the evaporation KY4 .
Thus, the non-perturbative solution (32) with (33) has no singularity. We will see in Sec.IV.3.2 that a quantum repulsive force is generated inside and plays a key role in the resolution of the singularity.
@@ IV.3 Semi-classical gravity condensate
where $K$ is a $\mathcal{O}(1)$ number to be fixed. This is an exponentially large redshift KMY . To see this effect, setting $r=R-\Delta r$ ($\Delta r\ll R$), (37) becomes $-g_{tt}^{*}(r)\sim e^{-\frac{R\Delta r}{\sigma\eta}}$. This indicates that due to the strong redshift, a deep region with $\Delta r\gg\mathcal{O}\left(\frac{nl_{p}^{2}}{a_{0}}\right)$ is almost frozen in the outside time $t$. As a result, the information carried by the excited quanta is kept inside for an exponentially long time KY2 , and the gravity condensate looks almost black. (We will discuss its phenomenological effect in Sec.VII.)
IV.3.1 Origin of (18) and formation process
We have started from the semi-classical static condition (18) and reached uniquely the entropy-maximizing metric (32). We here discuss the origin of (18).
First, as discussed around Fig.4, in order for the whole region to be static, there cannot exist a trapped region, and the condition (18) should hold as a physical condition that takes quantum effects into account. In fact, we can check explicitly that this is the case for specific static spherically symmetric metrics such as self-gravitating thermal radiation (see Appendix A).
Second, a possible origin of (18) is a local sufficient condition for the Bousso bound, Eq.(1.9) in Ref.FMW (given by (52) in our paper). We note that the entropy-density inequality (19) is equivalent to the condition (18) for highly-excited configurations in which the relation (11) holds. In Sec.VI.2.1, we will show that the inequality (19) is a spherical static version of (52). Therefore, the static condition (18) is valid for highly-excited spherical static configurations where (52) holds. Note that (52) is a hypothesis motivated by a phenomenological argument FMW , and that it is not clear yet which is more fundamental, (18) or (52).
Third, in order to explore a dynamical origin of (18), let us consider the formation process of the gravity condensate and investigate how the condition (18) is saturated as a result of the time evolution. We first note that the gravity condensate has, by construction, the maximum entropy and should be obtained by a reversible process in thermodynamics (since, in general, entropy is kept at the maximum for the external parameters given at each step of such processes). Suppose, then, that a collection of many spherical radiations at Hawking temperature comes together slowly due to self-gravity. See Fig.5.
@@ IV.3.1 Origin of (18) and formation process
Figure 5: Reversible formation of the semi-classical gravity condensate in a heat bath of Hawking temperature.
@@
This corresponds to a reversible formation process in a heat bath and can be realized operationally by slowly changing the temperature of the heat bath to match the Hawking temperature of the energy at each stage KY1 .
@@
We can use (1) and make a self-consistent model that describes this process KMY (see Appendix in Ref.KY5 for a short review). We first note that in this formation process, the metric changes in time, and that generically, particle creation occurs in a time-dependent spacetime BD ; Barcelo . As in Fig.5, we model the radiation as a collection of concentric spherical null shells with small energy $\frac{\Delta a_{i}}{2G}$ and describe the region between the $i$-th and $(i+1)$-th shells by the Vaidya metric with energy $\frac{a_{i}}{2G}$, corresponding to the Bondi energy inside the $i$-th shell with radius $r_{i}$ (here, $\Delta a_{i}\equiv a_{i}-a_{i-1}>0$ and $r_{i}<r_{i+1}$). We can solve the WKB-approximated time evolution of s-waves of $n_{s}$ scalar fields that come in the Minkowski vacuum state and propagate through the metric, to build a formula for the energy flux of the particles created around each shell ($n_{s}=\mathcal{O}(1)\gg 1$). Then, by solving (1) with the flux formula self-consistently, we can analyze the time evolution of both the shrinking shells and the geometry, [^15] It is often believed that the backreaction from evaporation is negligible when considering the formation process of a black hole with mass $\frac{a_{i}}{2G}$ since the timescale of the evaporation $\Delta t\sim\frac{a_{i}^{3}}{l_{p}^{2}}$ is much longer than the timescale of the collapse $\Delta\tau\sim a_{i}$. However, this naive expectation is wrong because $t$ and $\tau$ are the time coordinate at infinity and that of the comoving observer along the collapsing matter, respectively, and it does not make sense to compare such two timescales. In fact, one can examine the time evolution of the evaporation and collapse in a common time coordinate, and check that the both effects compete near $r=a_{i}$ and the evolution (38) occurs KMY ; KY4 . to see that around each shell, Hawking-like radiation occurs with temperature $\frac{\hbar}{4\pi a_{i}}$ and intensity $\sigma_{s}\sim n_{s}l_{p}^{2}$, and that each shell approaches just outside the shrinking Schwarzschild radius $a_{i}$ due to the evaporation,
$r_{i}\to a_{i}+\frac{2\sigma_{s}}{a_{i}},$ (38)
without forming a trapped region. Taking a continuum limit $\Delta a_{i}\to 0$, we can construct the interior metric (32) with $\sigma=\sigma_{s}$ and $\eta=1$, indicating that the gravity condensate is formed in this evolution.
From the evolution (38), we can see a dynamical saturation of (18). When the size of the gravity condensate is $\sim a_{i}$, the temperature of the bath is tuned at $T_{i}=\frac{\hbar}{4\pi a_{i}}$, and the radiation emitted from the bath comes to $r=a_{i}+\frac{2\sigma_{s}}{a_{i}}$ and becomes a constituent of that part of the gravity condensate. Then, the temperature is blueshifted to $\frac{T_{i}}{\sqrt{1-\frac{a_{i}}{r_{i}}}}\sim\frac{m_{p}}{\sqrt{n_{s}}}\sim\epsilon_{max}$ (from (38)), which means that the left-hand side of (18) is given by $\frac{\hbar}{\epsilon_{max}}\sim\sqrt{n_{s}}l_{p}$. Similarly, the right-hand side can be calculated as $\sqrt{r_{i}(r_{i}-a_{i})}\sim\sqrt{n_{s}}l_{p}$. Therefore, the condition (18) is saturated.
@@
Furthermore, this reversible formation is consistent with Bekenstein’s idea Bekenstein . At each stage, a photon at $T_{i}=\frac{\hbar}{4\pi a_{i}}$ comes to the gravity condensate of size $R_{i}\approx a_{i}+\frac{2\sigma_{s}}{a_{i}}$. Since the wavelength $\sim a_{i}$ and the size $R_{i}$ are almost the same, the probability for the photon to enter the condensate is roughly one-half, leading to 1 bit of information Bekenstein . From this and the previous paragaraph, the photon after the approaching (38) has the local energy $\epsilon_{max}$ and the 1 bit of information about its existence. Repeating this process, the uniform dense structure like Fig.2 appears, and the entropy follows the area law (25). In this sense, the gravity condensate realizes Bekenstein’s idea, including the interior description at the level of (1).
Thus, we can conclude from the above arguments that, at least as long as we apply this model by considering a spherical collapsing matter as a collection of many spherical shells with small energy, the condition (18) naturally holds as a result of the dynamics of (1). [^16] See also the 7th prospect in Sec.VII. Then, regarding the model as a realization of thermodynamically reversible formation, the condition (18) and hence the bound (21), is a consequence of the second law of thermodynamics at the level of the semi-classical Einstein equation (1). However, the model is not complete yet; for example, it only can reproduce the interior metric with $\eta=1$ because it does not incorporate the effects of interaction between constituent quanta KY1 ; KY4 . Therefore, further research will be needed to gain a deeper understanding of the origin of the condition (18), although it could be expected from the phenomenological but generic argument below Fig.4 that (18) should hold at least for highly-excited spherical static configurations.
IV.3.2 Origin of large tangential pressure $\langle T^{\theta}{}_{\theta}\rangle_{*}$
Before closing this section, we discuss the origin of the large tangential pressure $\langle T^{\theta}{}_{\theta}\rangle_{*}$ in (IV.3). Let us imagine that self-gravity brings together a large number of excited modes representing a collapsing matter, and the gravity condensate is formed (as in Fig.5). As the curvature increases, the vacuum fluctuations of the other modes with various angular momenta are induced to produce the quantum pressure as a 4D non-perturbative effect, even if the collapsing matter (such as dust) does not have a classical pressure KY4 (see Appendix B for a review). [^17] In the multi-shell model described above, we can use Israel’s junction condition and see that a strong surface pressure occurs on each shell due to the 4D energy-momentum conservation and the energy flux of the created particles KMY . It can be shown that its continuum version is $\langle T^{\theta}{}_{\theta}\rangle_{*}$ KY4 .
This can also be understood through the 4D Weyl anomaly, which is determined by the curvatures $\mathcal{R}$: $\langle T^{\mu}{}_{\mu}\rangle\sim\hbar n\mathcal{R}^{2}$ BD . From (IV.3), we have $\langle T^{\theta}{}_{\theta}\rangle_{*}\approx\frac{1}{2}\langle T^{\mu}{}_{\mu}\rangle_{*}$ as the leading relation for $r\gg l_{p}$ and obtain the large pressure from the large curvatures (34) KY3 .
Because of the pressure, the excited quanta are not concentrated in the center but are distributed throughout the interior, forming the gravity condensate. The curvature remains finite as in (34) self-consistently, the center has only small energy, and no singularity appears. This is the non-perturbative dynamical mechanism to resolve the singularity.
@@ V.1.1 Local temperature
We first find the local temperature $T_{loc}^{*}(r)$. We note that the excited quanta composing the part around $r$ of the gravity condensate are in radially accelerated motion against self-gravity to stay there as in Fig.2. The required acceleration is
@@
which is (from (33)) $\sim\epsilon_{max}$ consistent with (10). Thus, the excited quanta behave typically like a local thermal state in the radial direction at temperature (40).
This is kinematical and can be applied to any type of local degrees of freedom in the metric (32), due to the universality of the Unruh effect. In Sec.IV.3.1, on the other hand, we have got $T_{loc}(R_{i})\sim\epsilon_{max}$, which derives (40) dynamically through particle creation (see Ref.Y1 for its explicit derivation). Therefore, the kinematical and dynamical results coincide, and (40) is robust.
@@
Therefore, we can understand (40) as follows. At each stage of the adiabatic formation in Sec.IV.3.1, the local temperature of the outermost layer is $T_{loc}(r)\sim\epsilon_{max}$. After the subsequent radiation (or “shell”) covers this part, the local temperature is kept through the redshift. Only during $\Delta t<\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$ can the gravity condensate exist consistent with Tolman’s law. In this sense, the entropy-maximizing configuration is not in global equilibrium but in radially local one  [^18] This non-global equilibrium makes the difference from the result of Ref.Oppenheim , and is consistent with the self-consistent state $|\psi\rangle_{*}$ (mentioned above (33)). See Prospect 3 in Sec.VII for equilibrium states in quantum gravity.. This is consistent with the result that Tolman’s law does not lead to maximum entropy (see Appendix D). Furthermore, we will see in Sec.VI.2.2 that the saturation of a local Bousso bound is realized by the Unruh temperature, not by the Tolman’s one.
@@ V.2 Surface
We here determine the relation of the size $R$ and the total energy $M_{0}=\frac{a_{0}}{2G}$ from thermodynamics Y1 . Imagine that the gravity condensate is in equilibrium with a heat bath of temperature $T_{0}$ during $\Delta t<\mathcal{O}(e^{a_{0}^{2}/l_{p}^{2}})$. From thermodynamic relation $T_{0}dS=dM_{0}$ and (43), the equilibrium temperature is determined as the Hawking temperature GH :
@@ VI.1 Derivation of the Bousso bound
We show that the Bousso bound holds in our class of configurations. We first check that the entropy evaluated on a $t$-constant spacelike hypersurface $\Sigma_{t}$ inside the surface at $r=R$, which we have considered so far, agrees with one evaluated on a spherical ingoing null hypersurface $\Sigma_{L}$ starting from the surface at $r=R$ and converging at $r=0$ (a light sheet) Bousso1 . We suppose here that all configurations in our class satisfy (1) and are regular. [^21] At lease, we have seen in Sec.IV.2.2 that the entropy-maximized, most dense, configuration is regular. Applying $\nabla_{\mu}\tilde{s}^{\mu}=0$ to a finite spacetime region $V$ enclosed with $\Sigma_{t}$ and $\Sigma_{L}$ ($\partial V=\Sigma_{L}-\Sigma_{t}$ as in Fig.6)